# Can we accurately read or write quantum data?

### v2

@@ III Generalization to mixed states
$\operatorname{tr}_{S}\rho(0)=\Pi_{\varnothing}\big{(}\operatorname{tr}_{S}\rho(0)\big{)}\Pi_{\varnothing}^{\dagger},$ (13)
where $\operatorname{tr}_{S}\big{(}\rho(0)\big{)}$ is the reduced density operator obtained by partial trace over the observed system. The meaning of equation (13) is that the reduced density operator representing the pointer vanishes outside of the eigenspace representing the “ready” state of the pointer.
Condition (13) may be too relaxed, because it allows the observed system and the measuring device to be entangled from the start. If they are not entangled, $\rho(0)$ has the form $\rho(0)=\rho(0)_{S}\otimes\rho(0)_{M}$. Then already $\operatorname{tr}_{S}\rho(0)=\operatorname{tr}_{S}\big{(}\rho(0)_{S}\otimes\rho(0)_{M}\big{)}=\rho(0)_{M}$ and condition (13) becomes $\Pi_{\varnothing}\rho(0)_{M}\Pi_{\varnothing}^{\dagger}=\rho(0)_{M}$. Therefore, condition (13) is more relaxed than normally needed, but we will allow it for the sake of full generality.
@@
$\operatorname{tr}_{M}\rho(0)=\mathsf{P}_{\lambda}\big{(}\operatorname{tr}_{M}\rho(0)\big{)}\mathsf{P}_{\lambda}^{\dagger}$ (14)
at the end of the interaction the total state satisfies
$\rho(T)=\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}\rho(T)\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}^{\dagger}.$ (15)
@@
$\operatorname{tr}_{M}\rho(T)=\mathsf{P}_{\lambda}\big{(}\operatorname{tr}_{M}\rho(T)\big{)}\mathsf{P}_{\lambda}^{\dagger}.$ (16)
@@
$\operatorname{tr}_{S}\rho(t)=\Pi_{\lambda}\big{(}\operatorname{tr}_{S}\rho(t)\big{)}\Pi_{\lambda}^{\dagger}$ (17)
for any $\lambda$ and $t\in[T,T^{\prime}]$.
These conditions seem to give the most relaxed, the most general characterization of accurate measurements or preparations of mixed states for sharp observables. If we try to generalize it even more by considering unsharp observables (positive operator-valued measures [8, 6]) instead of the projective observable $\mathsf{A}$, measurements or preparations will automatically be unable to accurately distinguish states, by construction.
@@
$\big{(}\mathsf{B},\mathsf{C}\big{)}_{\textnormal{HS}}:=\operatorname{tr}\big{(}\mathsf{B}^{\dagger}\mathsf{C}\big{)}.$ (18)
@@
$\widetilde{\operatorname{tr}}_{S}(\rho)$ $:=\mathsf{I}_{S}\otimes\operatorname{tr}_{S}(\rho)$ (20)
$\widetilde{\operatorname{tr}}_{M}(\rho)$ $:=\operatorname{tr}_{M}(\rho)\otimes\mathsf{I}_{M}.$
They are linear operators, because the partial traces are linear maps. Conditions (13)–(17) can be rewritten in terms of the operators $\widetilde{\operatorname{tr}}_{S}$ and $\widetilde{\operatorname{tr}}_{M}$ and the projectors
$\widetilde{\mathsf{P}}_{\lambda}$ $:=\mathsf{P}_{\lambda}\otimes\mathsf{I}_{M},$ (21)
$\widetilde{\Pi}_{\lambda}$ $:=\mathsf{I}_{S}\otimes\Pi_{\lambda},$ $\widetilde{\Pi}_{\varnothing}$ $:=\mathsf{I}_{S}\otimes\Pi_{\varnothing},$
@@
In terms of the operators (20) and (21), condition (13) becomes $\widetilde{\operatorname{tr}}_{S}\rho(0)=\widetilde{\Pi}_{\varnothing}\big{(}\widetilde{\operatorname{tr}}_{S}\rho(0)\big{)}\widetilde{\Pi}_{\varnothing}^{\dagger}$. Since the partial trace $\operatorname{tr}_{S}$ takes density operators on $\mathcal{H}_{S}\otimes\Pi_{\varnothing}\mathcal{H}$ into density operators on $\Pi_{\varnothing}\mathcal{H}$, the operator $\widetilde{\operatorname{tr}}_{S}$ takes density operators on $\mathcal{H}_{S}\otimes\Pi_{\varnothing}\mathcal{H}$ into Hermitian operators on $\mathcal{H}_{S}\otimes\Pi_{\varnothing}\mathcal{H}$. Because of this, when condition (13) is satisfied,
@@
Taking into account that density operators on $\mathcal{H}$ are also vectors in the vector space $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$, we apply will Lemma 1 to the following case. As a Hilbert space we use the vector space of positive semi-definite trace-class operators with finite Hilbert-Schmidt norm. Let us denote it by $\widetilde{\mathcal{H}}$. It is a real Hilbert space. Anticipating the potential objection that Lemma 1 was proved for complex, not for real Hilbert spaces, we can take the complexified $\widetilde{\mathcal{H}}\otimes\mathbb{C}$ instead of $\widetilde{\mathcal{H}}$, and extend the Hilbert-Schmidt inner product (18) to a Hermitian scalar product, obtaining a Hilbert space. Then the unitary evolution operators $\mathbf{U}_{t}$, extended by linearity to $\widetilde{\mathcal{H}}\otimes\mathbb{C}$, act unitarily on $\widetilde{\mathcal{H}}\otimes\mathbb{C}$. As the vector $\Psi$ from Lemma 1 we will take $\rho(0)$, which is a vector from the space $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$. As a closed subspace $\mathcal{H}^{\prime}$ we take the subspace $\widetilde{\Pi}_{\lambda}\widetilde{\mathcal{H}}\otimes\mathbb{C}$.
@@
IV Discussion
@@
We have seen that if the Hamiltonian is bounded from below, accurate measurements and accurate preparations are impossible. Therefore, accurate reading and writing data, being measurements and preparations, are also impossible. This doesn’t prevent us from getting asymptotically closer to perfect accuracy, just from ever reaching it. This limitation is similar to the other limitations of quantum measurements reviewed in the Introduction, but if the Hamiltonian is bounded from below it is stronger, because it prevents measurement accuracy and non-disturbance independently and for all situations, and it does not need an additive conserved observable that does not commute with the observable $\mathsf{A}$.
@@
Given that all that we know about the world, even its macroscopic properties, we know through observations, which are ultimately quantum, this would make everything we know about the external world to be inaccurate. All we do in the world, even macroscopically, ultimately rely on quantum preparations, making our actions imprecise. But is this really so? We certainly have very complex electronic devices, including programmable classical computers, and they seem to work very well. They also seem scalable, at least so far we didn’t encounter any limitation of the memory they can have. While they are classical computers, they are still quantum systems, and not only because transistors employ quantum effects, but because the world is quantum. If the existence of classical computers contradicts our results, could it be because the Hamiltonian is actually unbounded? We don’t have a quantitative prediction of the possible accuracy in terms of the size or other properties of the measurement device, so it would be rushed to draw some conclusions from this. But we definitely need to explore this and find out.
@@
Only time will tell how to solve this dilemma. To avoid closing on a pessimistic note regarding accurate measurements and preparations, even if it turns out that they limit the possibility of scalable quantum computers, we have to remember what we learned from the history of quantum mechanics. Other results, initially perceived as limitations, are behind the incredible development of quantum technologies and other applications: results like complementarity, Bell’s theorem [2], the Kochen-Specker theorem [14], the no cloning theorem [29, 36] and so on, in the hands of visionary thinkers turned out to provide the revolutionary power of quantum technologies. Maybe these results rejected some of our initial intentions, but at the same time they opened unexpected avenues.
Acknowledgments. The author thanks Mohammad Hamed Mohammady for valuable comments and suggestions offered to a previous version of the manuscript. Nevertheless, the author bears full responsibility for the article.

### v3

@@ 4 Generalization to mixed states
[…]
@@
[…]
then, at the end of the interaction, the total state satisfies
$\rho(T)=\big(\mathsf{I}_{S}\otimes\Pi_{\lambda}\big)\rho(T)\big(\mathsf{I}_{S}\otimes\Pi_{\lambda}\big)^{\dagger}$ (13)
and conversely.
@@
[…]
@@
[…]
for any $t\in[T,T^{\prime}]$.
These conditions give the most relaxed, the most general characterization of accurate measurements or preparations of mixed states for sharp observables. If we try to generalize it even more by considering unsharp observables (positive operator-valued measures [29, 25]) instead of the projective observable $\mathsf{A}$, measurements or preparations will automatically be unable to accurately distinguish states, by construction.
@@
[…]
@@
[…]
$\widetilde{\tr}_{M}(\rho) :=tr$$=\tr_{M}(\rho)\otimes\mathsf{I}_{M}.$
[…]
$\widetilde{\mathsf{P}}_{\lambda}$ $=\mathsf{P}_{\lambda}\otimes\mathsf{I}_{M},$ (19)
$\widetilde{\Pi}_{\lambda}$ $=\mathsf{I}_{S}\otimes\Pi_{\lambda},$ $\widetilde{\Pi}_{\varnothing}$ $=\mathsf{I}_{S}\otimes\Pi_{\varnothing},$
@@
[…]
@@
Taking into account that density operators on $\mathcal{H}$ are also vectors in the vector space $\textnormal{Herm}\big(\mathcal{H}\big)$, we will apply Lemma 1 to the following case. As a Hilbert space we use the vector space of positive semi-definite trace-class operators with finite Hilbert-Schmidt norm. Let us denote it by $\widetilde{\mathcal{H}}$. It is a real Hilbert space. Anticipating the potential objection that Lemma 1 was proved for complex, not for real Hilbert spaces, we can take the complexified $\widetilde{\mathcal{H}}\otimes\mathbb{C}$ instead of $\widetilde{\mathcal{H}}$, and extend the Hilbert-Schmidt inner product (16) to a Hermitian scalar product, obtaining a complex Hilbert space. Then the unitary evolution operators $\mathbf{U}_{t}$, extended by linearity to $\widetilde{\mathcal{H}}\otimes\mathbb{C}$, act unitarily on $\widetilde{\mathcal{H}}\otimes\mathbb{C}$. As the vector $\Psi_{0}$ from Lemma 1 we will take $\rho(0)$, which is a vector from the space $\textnormal{Herm}\big(\mathcal{H}\big)$. As a closed subspace $\mathcal{H}^{\prime}$ we take the subspace $\widetilde{\Pi}_{\lambda}\widetilde{\mathcal{H}}\otimes\mathbb{C}$.
@@
5 Possible objections
Despite the generality of Theorem 1 and especially of Theorem 2, a mathematical result is as strong as its assumptions. Therefore, if one wants to reject these results to save both the Hamiltonian’s boundedness from below and the accuracy of quantum measurements and preparations, one can try to reject these assumptions.
Objection 1 (No perfect isolation).
Assumption 2 is not true, because the system composed of the measuring device and the observed system can never be perfectly isolated from the environment.
Reply.
We can never be sure that a fluctuation from the environment doesn’t happen, making the outcomes of the measurement accurate by a pure accident. However, if we worry that the measurement is not perfectly isolated, we can take as our system $M$ the whole laboratory. We can even assume that the laboratory is on a spaceship in an inertial motion, far away from any other system. Or we can take the whole universe as our system, if we can assume that the universe is an isolated system. Then, we replace the Hilbert space $\mathcal{H}_{M}$ with $\mathcal{H}_{M}\otimes\mathcal{H}_{E}$, and the pointer observable $\mathsf{Z}$ with $\mathsf{Z}\otimes\mathsf{I}_{E}$, where $\mathcal{H}_{E}$ represents the environment, and $\mathsf{I}_{E}$ is the identity operator on $\mathcal{H}_{M}$. Then, the proofs of Theorems 1 and  2 remain unchanged. ∎
Objection 2 (Time-dependence of the Hamiltonian).
Lemma 1 from Ref. [24] assumes a time-independent Hamiltonian; however, the Hamiltonian is, in fact, time dependent due to the presence of the interaction term $\mathsf{H}_{S+M}$ in $\mathsf{H}=\mathsf{H}_{S}\otimes\mathsf{I}_{M}+\mathsf{I}_{S}\otimes\mathsf{H}_{M}+\mu\mathsf{H}_{S+M}$, which turns on at time 0 and turns off at time $t=T$. To treat the problem properly, this time dependence must be taken into account, as it is the central effect under investigation. If one does so, it will be found that the conclusions of Lemma 1 do not hold.
Reply.
Consider, for example, the spin measurement of a Silver atom using the Stern-Gerlach device, the interaction is considered irrelevant before the Silver atom enters the magnetic field of the device and after it leaves it. This is why many discussions of the measurement process use a Hamiltonian of the form $\mathsf{H}=\mathsf{H}_{S}\otimes\mathsf{I}_{M}+\mathsf{I}_{S}\otimes\mathsf{H}_{M}+\mu\mathsf{H}_{S+M}$, with $\mu=0$ outside of the time interval $[0,T]$. In practice, however, there is no perfect way to ensure that no interactions occur outside of $[0,T]$.
However, our proofs don’t need to assume that the interaction is exactly zero before the measurement starts. The Silver atom could very well already feel the magnetic field of the Stern-Gerlach device at $t=0$ or before, and yet, the total Hamiltonian can be time-independent, ensuring the validity of equation (3). We don’t even need to assume that it is exactly zero for $t<0$ and $t>T$, because all we used from Assumption 3 is that the pointer state is an eigenstate of the pointer observable, at least for a finite time interval $T,T^{\prime}]$, as in equation (7) or, more generally, as in equation (15). The Hamiltonian can be assumed to be time-independent due to Assumption 2, as explained in the Reply to Objection 1.
Possible external influences that can make the Hamiltonian time-dependent are either noise from the environment, or controlled interactions that aim to restore the accuracy. In the first case, it is unlikely that external noise will restore the accuracy of the measurement, it will rather introduce more error. In the second case, intentional interactions can restore the accuracy of the measurement only if controlled by someone who already accurately knows the state for all the systems involved. ∎
Objection 3 (Challenging pointer persistence).
Assumption 3 is, well, an unproven assumption. Such constructs have no place in a formal proof.
Reply.
There are two points to unpack in this objection. First, even if one thinks that Assumption 3 is unproven, it still has a place in a formal proof. Every formal proof should transparently state the assumptions under which the result is derived, so that the domain of applicability of the result can be known. Then, if one of the assumptions is rejected, the proof does not necessarily apply in that case. What has no place in formal proof is not having assumptions, but having hidden assumptions.
So if one thinks that Assumption 3 is wrong, and that we can do measurements even if their results are not temporarily recorded, one is free to reject the conclusion of Theorem 1. But one can’t reject the proof itself on this basis alone, only the conclusion.
However, rejecting Assumption 3 will not save the accuracy, quite the contrary:
A perfectly accurate result that is not recorded can’t be seen by the experimentalist or anyone else, so it is totally useless.
If Assumption 3 is not made, the experimentalist can’t trust the result, since it may be perturbed before being read. This would defeat the whole purpose of Objection 3, by making the result inaccurate. ∎
Objection 4.
The proof of Theorem 1 is based on the assumption of unitarity. But at the end the Projection Postulate is applied to reduce the superposition from equation (3) to a single term. So unitarity is violated at $t=T$, and Lemma 1 can’t be applied.
Reply.
The proof derives a contradiction from the cases when no projection is needed, because the observed system was in an eigenstate of the observable. See the details of the proof, Remark 1, and Section §3. ∎
Question 1.
The proof of Theorem 1 is based on the condition (4), respectively (5), and so it can be understood as a proof by reductio-ad-absurdum that these conditions can’t be satisfied independently of the initial state of the observed system. And, indeed, there are situations when a separable state evolves into a separable state [30].
Answer.
Indeed, one can take the proof of Theorem 1 as a proof by reductio-ad-absurdum that conditions (4) and (5) can’t be satisfied independently of the initial state of the observed system. That is, it is not excluded that the pointer state is not in the “ready” state $\varphi_{\varnothing}(0)$ for all possible initial states, but if the observed system is already in an eigenstate $\psi_{\lambda}$ of the observable, the pointer had to be in an eigenstate for $\lambda$, and not of $\varphi_{\varnothing}(0)$ as normally assumed. This will allow accuracy even if the Hamiltonian is bounded from below. It will also have the advantage of ensuring a single outcome without breaking the unitary evolution with a projection or collapse. Such a single-world unitary (no-collapse) proposal was shown in [31] to require apparently very fine-tuned initial conditions that depend on the future measurement. In this case, both the observed system and the pointer state have to be, at $t=0$, eigenstates corresponding to the outcome $\lambda$. This appears as “superdeterministic” or “retrocausal”, although it can also be interpreted as the result of a global consistency condition in the four-dimensional relativistic block universe [32, 33]. In [34, 33] it was shown that this is the only way to ensure the conservation laws (Everett’s interpretation also ensures them, but for the totality of the branches, while in each branch conservation laws are violated). In [35], a new version of this single-world unitary proposal, which separates the observables for their physical meaning to eliminate the apparent superdeterminism or retrocausality, was proposed. While, in the current context, it seems that rejecting conditions (4) and (5) can allow for accuracy even if the Hamiltonian is bounded from below, it does not present a clear mechanism how this would be possible. By contrast, since unitarity without projection is a very strong constraint, it may restrict the accuracy even more. But the possibility from Question 1 is indeed open. ∎
6 Discussion
@@
We have seen that if the Hamiltonian is bounded from below, accurate measurements and accurate preparations are impossible. Therefore, if this is true, accurate reading and writing data, being measurements and preparations, are also impossible. This doesn’t prevent us from getting asymptotically closer to perfect accuracy, just from ever reaching it. This limitation is similar to the other limitations of quantum measurements reviewed in the Introduction, but if the Hamiltonian is bounded from below it is stronger, in some sense, than the Wigner-Araki-Yanase theorem [2, 3, 4], because it prevents measurement accuracy and non-disturbance independently and for all situations, and it does not need to use an additive conserved observable that does not commute with the observable $\mathsf{A}$.
@@
Given that all that we know about the world, even its macroscopic properties, we know through observations, which are ultimately quantum, this would make everything we know about the external world to be inaccurate. All we do in the world, even macroscopically, ultimately rely on quantum preparations, making our actions imprecise. But is this really so? We certainly have very complex electronic devices, including programmable classical computers, and they seem to work very well. They also seem scalable, at least so far we didn’t encounter any limitation of the memory they can have. While they are classical computers, they are also quantum systems, and not only because transistors employ quantum effects, but because the world is quantum. So we can consider the widespread and ubiquitous usage of information technology in our everyday life as the largest crowdsourced experiment that tested the accuracy of reading and writing bits. In fact, this experiment is a natural continuation of the longest crowdsourced experiment that showed that we can keep records of past events, for example by writing them down on a piece of paper. Even a mark on a tally stick is a physical state that can be distinguished from other physical states by observations that are, ultimately, quantum. If the existence of recording devices or classical computers contradicts our results, could it be because the Hamiltonian is actually unbounded? We don’t have a quantitative prediction of the possible accuracy in terms of the size or other properties of the measurement device, so it would be rushed to draw some conclusions from this. But we definitely need to explore this and find out.
@@
Only time will tell how to solve this dilemma. To avoid closing on a pessimistic note regarding accurate measurements and preparations, even if it turns out that they limit the possibility of scalable quantum computers, we have to remember what we learned from the history of quantum mechanics. Other results, initially perceived as limitations, are behind the incredible development of quantum technologies and other applications: results like complementarity, Bell’s theorem [41], the Kochen-Specker theorem [42], the no cloning theorem [43, 44] and so on, in the hands of visionary thinkers turned out to provide the revolutionary power of quantum technologies. Maybe these results rejected some of our initial expectations and intentions, but at the same time they opened unexpected avenues.
Acknowledgments. The author thanks Mohammad Hamed Mohammady and anonymous referees for valuable comments and suggestions offered to a previous version of the manuscript. Nevertheless, the author bears full responsibility for the article.